OpenStudy (anonymous):

hey guys hope your well ! so I have this evaluating proof problem ,can anyone help me plzzz. Its geometric series , question is : suppose m and k are integers with m

OpenStudy (anonymous):

\[\sum_{n=m}^{k} ar ^{n}\]

OpenStudy (anonymous):

\[S(n) = \frac{ a _{1}(1-r) }{ 1-r }\]

OpenStudy (anonymous):

i know this formula will drive the same way this formula did ^ but just dont know how to set up ?

OpenStudy (anonymous):

@IrishBoy123

OpenStudy (anonymous):

@IrishBoy123 mind helping?

OpenStudy (irishboy123):

for first n terms you have \[\large \Sigma_{k=0}^{n-1} \ a \ r^k = a \frac{1-r^n}{1-r}\]

OpenStudy (irishboy123):

so you can calculate 2 summations, each starting at zero and ending somewhere that leaves you with the interval you want to sum over a bit like S(8 to 10) = S(10) - S(7)

OpenStudy (irishboy123):

so methodically plug into that formula and see where you go

OpenStudy (anonymous):

ok thank you

OpenStudy (irishboy123):

ouch, you are using k and m in your problem so i have removed them to make it cleaner: \[\large \Sigma_{i=0}^{n-1} \ a \ r^i = a \frac{1-r^n}{1-r}\]

OpenStudy (irishboy123):

i mean......i have replaced the k by i